# Transit Campus, Government ITI, Berhampur 760010, Odisha, India

### v1

@@ 1 Introduction
Numerous formation mechanisms for PBHs have been explored, including collapse from density fluctuations during different cosmic eras Carr:1974nx ; Grillo:1980rt ; Khlopov:1980mg ; Khlopov:1985fch ; Carr:1994ar ; Chakraborty:2022mwu , collapse in single and multi-field inflationary models Carr:1993aq ; Bullock:1996at ; Saito:2008em ; Kawai:2021edk ; Kawai:2021bye ; Choudhury:2023hvf ; Randall:1995dj ; Garcia-Bellido:2016dkw ; Braglia:2020eai ; Kawai:2022emp , collapse of cosmic string loops Hawking:1987bn ; Borah:2023iqo , and collapse during a First Order Phase Transition (FOPT) Crawford:1982yz ; Kawana:2021tde ; Baker:2021nyl ; Huang:2022him ; Kawana:2022olo ; Liu:2021svg ; Gouttenoire:2023naa ; Lewicki:2023ioy . Many of these processes have also been considered as potential sources of GWs Matarrese:1993zf ; Matarrese:1996pp ; Matarrese:1997ay ; Khlebnikov:1997di ; Easther:2006vd ; Easther:2007vj ; Choudhury:2013woa ; Binetruy:1981mr ; Vachaspati:1984gt ; Hindmarsh:1994re ; Witten:1984rs ; Hogan:1986dsh . We will focus on the FOPT origin of PBHs and GWs within the framework of the $U(1)_{B-L}$ extension of the Inert Doublet Model (IDM), where the FOPT is associated with the high-scale breaking of the $U(1)_{B-L}$ symmetry as well as electroweak symmetry breaking.
As a simple inert doublet extension of the Standard Model (SM) IDM was proposed to address the dark matter issue under the WIMP paradigm. It was soon realised that with the introduction of right handed neutrinos this model can explain the active neutrino masses radiatively Deshpande:1977rw ; Ma:2006fn ; Ma:2006km ; Barbieri:2006dq ; LopezHonorez:2006gr . However, the discovery of Higgs boson at 125 GeV and the ever-increasing precision of direct detection experiments excluding so much of the parameter space takes away the charm of even the neutrino augmented IDM. To address this we consider an additional $U(1)_{B-L}$ symmetry which can allow a significant first-order phase transition (FOPT) that can generate PBHs and GWs. As a consequence these PBHs can play the role of DM and the burden of satisfying the DM constraints by the inert scalar is lifted. Additionally this allows us a complementary probe of the scenario by means of GW astronomy.
@@ 3 First Order Phase Transition
which suggests that a larger value of $\beta/H$ denotes a faster FOPT. These are the important parameters of a FOPT which enters and takes pivotal participation in the calculation of the GW amplitude and frequency as well as the mass and abundance of the PBH population which we discuss next.
@@ 3.1 Creation of Gravitational Waves
Gravitational wave background of stochastic nature can be created from a FOPT in a few different ways, i.e. (i) collision of bubbles, (ii) sound waves in the plasma, and (iii) magnetohydrodynamic turbulence in the background plasma. The first two among these processes has been modeled mathematically with a degree of success, however, the third one is yet to be modeled to a comfortable degree of accuracy. For our cases, since the FOPT is strong enough, we consider only the collision of bubble walls for our cases. The energy density of the GW due to the bubble collision can be expressed as Huber:2008hg ; Caprini:2018mtu ,
@@
$\Omega_{\mathrm{bub}}(f)h^{2}=1.67\times 10^{-5}\left(\dfrac{100}{g_{*}}\right)^{1/3}\left(\dfrac{H}{\beta}\right)^{2}\left(\dfrac{\kappa_{\phi}\alpha}{1+\alpha}\right)^{2}\dfrac{0.11v_{w}^{3}}{0.42+v_{w}^{2}}\dfrac{3.8(f/f_{\mathrm{peak}}^{\mathrm{bub}})^{2.8}}{1+2.8(f/f_{\mathrm{peak}}^{\mathrm{bub}})^{3.8}},$ (14)
where $\kappa_{\phi}$ is the fraction of the latent heat energy that gets converted into the energy gradient of the field driving the FOPT, $v_{w}$ is the velocity of the bubble wall, and $f_{\mathrm{peak}}^{\mathrm{bub}}$ is the peak of the GW spectrum and can be expressed as,
@@
$f_{\mathrm{peak}}^{\mathrm{bub}}=1.65\times 10^{-5}\mathrm{~{}Hz}\left(\dfrac{g_{*}}{100}\right)^{1/6}\left(\dfrac{T_{p}}{100\mathrm{~{}GeV}}\right)\dfrac{0.62}{1.8-0.1v_{w}+v_{w}^{2}}\left(\dfrac{\beta}{H}\right).$ (15)
We use these expressions for our analysis of the GW spectrum coming from the strong FOPT in the underlying model.
@@ 3.2 Creation of Primordial Black Holes
Now we discuss the creation of PBHs from a FOPT. The basic mechanism is that since the nucleation of true vacuum bubble is probabilistic process, there may exist some hubble region in which bubble nucleation has not taken place, i.e., that region is still in false vacuum and dominated by vacuum energy whereas its surrounding regions are in true vacuum and dominated by radiation. As a result the expansion of the universe dilutes the surrounding energy density keeping the energy density of the concerned Hubble region constant resulting in a overdensity. If and when this overdensity crosses a certain threshold value, the region collapses into a PBH.
In order to quantitatively determine the properties of the PBH population such as the abundance, and the spin of PBHs as functions of the FOPT parameters, such as $\alpha$ and $\beta/H$, we follow the prescription of Ref. Banerjee:2023qya . The curvature perturbation due to a FOPT can be expressed as Liu:2022lvz ,
$\mathcal{P}_{\mathcal{R}}(k)=34.5[\sigma_{H}(\alpha,\beta/H)]^{2}(k\mathcal{H})^{3},$ (16)
where $\mathcal{H}(=aH)$ is the comoving Hubble parameter and $\sigma_{H}(\alpha,\beta/H)$ is a function which determines the amplitude of the perturbation, i.e., it increases (decreases) with $\alpha$ ($\beta/H$). The spectral moments can be expressed as Young:2019yug ,
$\sigma^{2}$ $=\dfrac{16}{81}\int\dfrac{dk}{k}\mathcal{P_{R}}(k)\left(\dfrac{k}{\mathcal{H}}\right)^{4}W^{2}(k,r_{m})T^{2}(k,r_{m}),$ (17a)
$\mu^{2}$ $=\dfrac{16}{81}\int^{\infty}_{0}\dfrac{dk}{k}(kr_{m})^{4}W^{2}(k,r_{m})T^{2}(k,r_{m})\mathcal{P}_{\mathcal{R}}(k)k^{2}.$ (17b)
where $\sigma$ and $\mu$ are the zeroth and first spectral moments respectively. The window ($W(k,r_{m})$) and the transfer ($T(k,r_{m})$) functions can be expressed as Young:2019yug ,
$W(k,r_{m})$ $=3\dfrac{\sin(kr_{m})-kr_{m}\cos(kr_{m})}{(kr_{m})^{3}},$ (18a)
$T(k,r_{m})$ $=3\dfrac{\sin(kr_{m}/\sqrt{3})-(kr_{m}/\sqrt{3})\cos(kr_{m}/\sqrt{3})}{(kr_{m}/\sqrt{3})^{3}}.$ (18b)
It is to be noted here that the top-hat window function and the transfer functions are used to smoothen the curvature perturbation in the scales we are interested in. Furthermore, here the threshold value of the gaussian component ($\delta_{l}$) of the density perturbation ($\delta$) is $\delta_{c,l-}=0.741$ which corresponds to the threshold of $\delta$, i.e., $\delta_{c}=0.535$ Banerjee:2023qya . The fraction of the universe that collapses into PBHs during the time of formation can be obtained as Young:2019yug ,
[…]
where $\nu=\delta_{l}/\sigma$, $\nu_{\mathrm{th}}=\delta_{c,l-}/\sigma$ and $\mathcal{K}$ is a numerical prefactor with depends on the threshold; we consider $\mathcal{K}=5.5$ throughout the article for $\delta_{c,l-}=0.741$ Musco:2020jjb ; Banerjee:2023qya . From this the density of the PBHs today can be obtained as Young:2019yug ,
$\Omega_{\mathrm{PBH}}=\int^{M_{\mathrm{max}}}_{M_{\mathrm{min}}}\mathrm{d}(\mathrm{ln}M_{H})\left(\dfrac{M_{\mathrm{eq}}}{M_{H}}\right)^{1/2}\beta_{0},$ (20)
where $M_{\mathrm{max}}$ and $M_{\mathrm{min}}$ are the maximum and the minimum horizon masses at which the PBHs form, and $M_{\mathrm{eq}}$ is the horizon mass at the matter-radiation equality. The final abundance can be calculated as,
$f(M_{\mathrm{PBH}})=\dfrac{1}{\Omega_{\mathrm{CDM}}}\dfrac{\mathrm{d}\Omega_{\mathrm{PBH}}}{\mathrm{d}(\mathrm{ln}M_{\mathrm{PBH}})},$ (21)
where $\Omega_{\mathrm{CDM}}$ is the DM density today. For simplicity we consider the mass of a PBH is to be the same as that of the Hubble region. Since the mass of the Hubble region depends only on the temperature of the universe, the mass of the PBH can be expressed as,
$M_{\mathrm{PBH}}=9.23\times 10^{31}\left(\dfrac{T}{\mathrm{GeV}}\right)^{-2}\mathrm{~{}g},$ (22)
where $T$ is the temperature of the universe during the FOPT. In general we identify this as the percolation temperature.
It is to be noted that, qualitatively, as $\alpha$ increases, the abundance of the PBH population increases simply because higher $\alpha$ signifies greater overdensity. However, the abundance is much more sensitive to the $\beta/H$ and it deceases significantly with increasing $\beta/H$. This can be explained from the idea that $\beta/H$ is equivalent to the inverse of the duration of the FOPT and higher $\beta/H$ leads to faster FOPT which does not allow the overdensity to grow high enough to collapse. According to Ref. Banerjee:2023qya , irrespective of how high $\alpha$ is, if $\beta/H\geq 27$, the abundance of PBH is vanishingly small and hence can be neglected. Furthermore, the spin of the PBH in this is case is also $\mathcal{O}(10^{-5})$ therefore we neglect them in our study.
@@ 4.2 Parameter Scan
We can further divide the parameter scan effectively into two parts, namely, (i) $\chi$-related parameters and (ii) $\Phi_{1,2}$-related parameters.
@@ 4.2.1 $\chi$-Related Parameters
In this part, we discuss the parameter ranges relevant for our work to understand the properties of the first FOPT, i.e. the one driven by the scalar field $\chi$. The parameters in question for this case are $g_{B-L}$, $y_{1,2,3}$, and $v_{\chi}$. As mentioned before, these parameters control the strength ($\alpha$), the inverse of the duration ($\beta/H$) and the relevant temperatures ($T_{p},~{}T_{n},~{}T_{c}$) of the FOPT, and those in turn control the PBH mass and abundance and GW spectrum. At first in Fig. 1 we show the dependence of $\alpha$ and $\beta/H$ on the couplings $g_{B-L}$ for two values of the Yukawa couplings, i.e., $y_{1,2,3}=0.1,~{}0.2$ while keeping the value of the VEV of the $\chi$ field fixed at $v_{\chi}=10^{8}\mathrm{~{}GeV}$.
@@
Figure 1: Dependence of (left) $\alpha$ and (right) $\beta/H$ values on $g_{B-L}$ for $y_{1,2,3}=0.1,~{}0.2$. In this case $v_{\chi}=10^{8}\mathrm{~{}GeV}$.
@@
It can bee seen from Fig. 1 that for a fixed value of $y_{1,2,3}$, $\alpha$ ($\beta/H$) increases (decreases) significantly as $g_{B-L}$ decreases. On the other hand lower value of the Yukawa coupling leads to higher value of $\alpha$. For $\beta/H$, this dependence is not as simple, i.e., for $g_{B-L}\lesssim 0.195$ higher yukawa coupling leads to higher $\beta/H$ and vice versa. It is also to be noted that both $\alpha$ and $\beta/H$ depend very weakly on $y_{1,2,3}$ as compared to $g_{B-L}$. This behaviour is also true for the dependence $\alpha$ and $\beta/H$ on the $v_{\chi}$. We illustrate this further in Fig. 2 where we show the dependence of $\alpha$ and $\beta/H$ on $g_{B-L}$ for different values of VEV, i.e., $v_{\chi}=10^{7}\mathrm{~{}GeV},~{}10^{8}\mathrm{~{}GeV}$ for constant Yukawa coupling $y_{1,2,3}=0.2$.
@@
Figure 2: Dependence of (left) $\alpha$ and (right) $\beta/H$ on the coupling $g_{B-L}$ for $v_{\chi}=10^{7}\mathrm{~{}GeV},~{}10^{8}\mathrm{~{}GeV}$. In both the cases $y_{1,2,3}=0.2$.
@@
It can be seen from the figure that for higher value of VEV, $\alpha$ is higher. However, similar to the previous case, for $g_{B-L}\lesssim 0.117$, the $\beta/H$ is higher for higher VEV and vice versa.
@@
Apart from $\alpha$ and $\beta/H$, the temperatures related to the FOPT, i.e., $T_{p}$, $T_{n}$, and $T_{c}$ are also important since the peak of the SGWB and the mass of the PBH depends heavily on them. We found that for fixed VEV (Yukawa coupling), all three temperatures have similar behaviour with varying gauge and Yukawa coupling (VEV). Therefore in Fig. 3 we only show the $T_{n}$ dependence.
Figure 3: Dependence of $T_{n}$ on $g_{B-L}$ (left) and $y_{1,2,3}$ while keeping $v_{\chi}=10^{8}\mathrm{~{}GeV}$ (right) and $v_{\chi}$ keeping $y_{1,2,3}=0.2$. The white shaded regions are the once where we do not get any FOPT.
@@
From the figure, it can be seen that at fixed VEV and Yukawa coupling, $T_{n}$ increases with increasing $g_{B-L}$. In this case as well, $T_{n}$ has very weak dependence on the Yukawa couplings. However, for fixed gauge and Yukawa couplings, $T_{n}$ increases with VEV.
Since the dependence of the PBH abundance is very sensitive to the FOPT parameters, especially $\beta/H$, in Fig. 4 we show the dependence of the $f_{\mathrm{PBH}}$ on $\beta/H$ for different values of $\alpha$.
Figure 4: The dependence of the PBH abundance on $\beta/H$ with varying $\alpha$. In this case the percolation temperature, $T_{p}=10^{6}\mathrm{~{}GeV}$.
From the figure, it is evident that at low $\alpha$ values, the abundance decreases with $\beta/H$ at a much slower rate than at high $\alpha$ values. Furthermore, any kind of realistic abundance only occurs between the $\beta/H$ values of 25.3 to 26.8 and there is a very small parameter range corresponding $g_{B-L}$ and $y_{1,2,3}$ which gives rise to $\beta/H$ values adequate for creation of PBHs with enough abundance. We show this in more details as follows.

### v2

@@ 1 Introduction
Numerous formation mechanisms for PBHs have been explored, including collapse from density fluctuations during different cosmic eras Carr:1974nx ; Grillo:1980rt ; Khlopov:1980mg ; Khlopov:1985fch ; Carr:1994ar ; Chakraborty:2022mwu , collapse in single and multi-field inflationary models Carr:1993aq ; Bullock:1996at ; Saito:2008em ; Kawai:2021edk ; Kawai:2021bye ; Choudhury:2023hvf ; Randall:1995dj ; Garcia-Bellido:2016dkw ; Braglia:2020eai ; Kawai:2022emp ; Heydari:2021gea ; Heydari:2021qsr , collapse of cosmic string loops Hawking:1987bn ; Borah:2023iqo , and collapse during a First Order Phase Transition (FOPT) Crawford:1982yz ; Kawana:2021tde ; Baker:2021nyl ; Huang:2022him ; Kawana:2022olo ; Liu:2021svg ; Gouttenoire:2023naa ; Lewicki:2023ioy ; Roshan:2024qnv ; Kanemura:2024pae ; Goncalves:2024vkj ; Banerjee:2023qya ; Lewicki:2024ghw . Many of these processes have also been considered as potential sources of GWs Matarrese:1993zf ; Matarrese:1996pp ; Matarrese:1997ay ; Khlebnikov:1997di ; Easther:2006vd ; Easther:2007vj ; Choudhury:2013woa ; Binetruy:1981mr ; Vachaspati:1984gt ; Hindmarsh:1994re ; Witten:1984rs ; Hogan:1986dsh . We will focus on the FOPT origin of PBHs and GWs within the framework of the $U(1)_{B-L}$ extension of the Inert Doublet Model (IDM), where the FOPT is associated with the high-scale breaking of the $U(1)_{B-L}$ symmetry as well as electroweak symmetry breaking.
As a simple inert doublet extension of the Standard Model (SM) IDM was proposed to address the dark matter issue under the WIMP paradigm. It was soon realised that with the introduction of right handed neutrinos this model can explain the active neutrino masses radiatively Deshpande:1977rw ; Ma:2006fn ; Ma:2006km ; Barbieri:2006dq ; LopezHonorez:2006gr . However, the discovery of Higgs boson at 125 GeV and the ever-increasing precision of direct detection experiments excluding so much of the parameter space takes away the charm of even the neutrino augmented IDM. To address this we consider an additional $U(1)_{B-L}$ symmetry which can allow a significant first-order phase transition (FOPT) that can generate PBHs and GWs Baldes:2023rqv ; Salvio:2023ynn . As a consequence these PBHs can play the role of DM and the burden of satisfying the DM constraints by the inert scalar is lifted. Additionally this allows us a complementary probe of the scenario by means of GW astronomy.
@@ 3 First Order Phase Transition
which suggests that a larger value of $\beta/H$ denotes a faster FOPT. Another quantity of interest is the reheating temperature, i.e., the temperature where the vacuum energy is converted into radiation and this can be expressed as,
$T_{\mathrm{reh}}=T_{p}(1+\alpha)^{1/4}.$ (14)
These are the important parameters of a FOPT which enters and takes pivotal participation in the calculation of the GW amplitude and frequency as well as the mass and abundance of the PBH population which we discuss next.
@@ 3.1 Creation of Gravitational Waves
Gravitational wave background of stochastic nature can be created from a FOPT in a few different ways, i.e. (i) collision of bubble walls or relativistic fluid shells, (ii) sound waves in the plasma, and (iii) magnetohydrodynamic turbulence in the background plasma. The first two among these processes has been modeled mathematically with a degree of success, however, the third one is yet to be modelled to a comfortable degree of accuracy. Furthermore, for extremely slow and supercooled FOPTs, there is another source of SGWB, which is the large curvature perturbations (which also creates the PBHs). For our cases, as it is shown in the subsequent sections of this article, there is one supercooled strong FOPT and there is one weaker FOPT driven by a polynomial potential. For the cases of the former FOPT, main contribution for GW will come from collision and curvature perturbations whereas for the later it will come from the sound waves. The brief description for both of these contributions are given as follows.
Collision and Curvature Perturbation
The GW in case of a strongly supercooled and slow FOPT can be expressed as Lewicki:2024ghw ,
@@
$\Omega_{\mathrm{GW}}^{\mathrm{strong}}h^{2}=1.6\times 10^{-5}\left(\dfrac{g_{*}}{100}\right)\left(\dfrac{g_{*s}}{100}\right)^{-4/3}\left(\Omega_{\mathrm{coll}}+\Omega_{\mathrm{curv}}\right),$ (15)
where $g_{*}$ ($g_{*s}$) is the relativistic (entropy) degrees of freedom at the time of creation of the GW, and $\Omega_{\mathrm{coll}}$ and $\Omega_{\mathrm{curv}}$ are the GW spectrum due to the collision of bubble walls/relativistic fluid shells and the curvature perturbation respectively.
The spectrum for the collision can be expressed as Lewicki:2022pdb ,
$\Omega_{\mathrm{coll}}=\left(\dfrac{\beta}{H}\right)^{-2}\dfrac{A(a+b)^{c}S_{H}(k,k_{\mathrm{max}})}{\left(b(k/k_{p})^{-a/c}+a(k/k_{p})^{b/c}\right)^{c}},$ (16)
where $k_{p}\approx 0.7k_{\mathrm{max}}\beta/H$ is the peak wavenumber, $A=5.1\times 10^{-2}$ is the amplitude of the spectrum, $a=b=2.4$ and $c=4$ Lewicki:2022pdb . Furthermore,
$k_{\mathrm{max}}$ $=1.6\times 10^{-7}\left(\dfrac{g_{*}}{100}\right)^{1/2}\left(\dfrac{100}{g_{*s}}\right)^{1/3}\dfrac{T_{\mathrm{reh}}}{\mathrm{GeV}}~{}~{}~{}\mathrm{Hz}$ (17)
$S_{H}(k,k_{\mathrm{max}})$ $=\left(1+\dfrac{\Omega_{\mathrm{CT}}(k_{\mathrm{max}})}{\Omega_{\mathrm{CT}}(k)}\left(\dfrac{k}{k_{\mathrm{max}}}\right)^{a}\right)^{-1},$ (18)
where $\Omega_{\mathrm{CT}}$ ($\propto k^{3}$ in radiation dominated universe) signifies the tail due to the causality Lewicki:2022pdb .
On the other hand, the spectrum due to the curvature perturbation can be expressed as Kohri:2018awv ; Espinosa:2018eve ; Inomata:2019yww ,
[…]
where $\mathcal{P}_{\zeta}$ is the curvature perturbation created due to the supercooled slow FOPT and $\mathcal{I}_{t,s}$ signifies the transfer function and can be expressed as,
$\mathcal{I}^{2}_{t,s}$ $=\dfrac{288(s^{2}+t^{2}-6)}{(t^{2}-s^{2})^{6}}$
×$\times\left(\dfrac{\pi^{2}}{4}(s^{2}+t^{2}-6)^{2}\theta(t^{2}-3)+\left(t^{2}-s^{2}-\dfrac{1}{2}(s^{2}+t^{2}-6)\ln\left|\dfrac{t^{2}-3}{3-s^{2}}\right|\right)^{2}\right).$ (20)
Sound Waves
The SGWB spectra due to sound waves can be expressed as Hindmarsh:2013xza ; Hindmarsh:2015qta ; Hindmarsh:2017gnf ,
@@
$\Omega_{\mathrm{sw}}h^{2}=4.13\times 10^{-7}(R_{*}H_{*})\left(1-\dfrac{1}{\sqrt{1+2\tau_{\mathrm{sw}}H_{*}}}\right)\left(\dfrac{\kappa_{\mathrm{sw}}\alpha}{1+\alpha}\right)^{2}\left(\dfrac{100}{g_{*}}\right)^{1/3}S_{\mathrm{sw}}(f),$ (21)
where, $S_{\mathrm{sw}}$ contains the information regarding the frequency dependence of the SGWB spectrum and can be expressed as,
$S_{\mathrm{sw}}(f)=\left(\dfrac{f}{f_{\mathrm{sw}}}\right)^{3}\left(\dfrac{4}{7}+\dfrac{3}{7}\left(\dfrac{f}{f_{\mathrm{sw}}}\right)^{2}\right)^{-7/2},$ (22)
the peak frequency of the spectrum is given by,
$f_{\mathrm{sw}}=2.6\times 10^{-5}(R_{*}H_{*})^{-1}\left(\dfrac{T_{\mathrm{reh}}}{100~{}\mathrm{GeV}}\right)\left(\dfrac{g_{*}}{100}\right)^{1/6},$ (23)
and the other relevant quantities are defined as,
$\tau_{\mathrm{sw}}H_{*}$ $=\dfrac{R_{*}H_{*}}{U_{f}},$ (24)
$U_{f}$ $=\sqrt{\dfrac{3}{4}\dfrac{\alpha}{1+\alpha}\kappa_{\mathrm{sw}}},$ (25)
$R_{*}H_{*}$ $=(8\pi)^{1/3}\mathrm{max}(v_{w},~{}c_{s})\left(\dfrac{\beta}{H}\right)^{-1}.$ (26)
Here $c_{s}=1/\sqrt{3}$ is the speed of sound in the fluid and $v_{w}$ is the bubble wall velocity for which we use the approximate form Lewicki:2021pgr ,
$v_{w}=\begin{dcases}\sqrt{\dfrac{\Delta V_{\mathrm{tot}}}{\alpha\rho_{\mathrm{rad}}}},&\sqrt{\dfrac{\Delta V_{\mathrm{tot}}}{\alpha\rho_{\mathrm{rad}}}}<v_{J},\\ 1,&\sqrt{\dfrac{\Delta V_{\mathrm{tot}}}{\alpha\rho_{\mathrm{rad}}}}\geq v_{J},\\ \end{dcases}$ (27)
where, $v_{J}$ is the Chapman-Jouget velocity and is given by Espinosa:2010hh ,
$v_{J}=\dfrac{1}{\sqrt{3}}\dfrac{1+\sqrt{3\alpha^{2}+2\alpha}}{1+\alpha}.$ (28)
Finally, for the sound wave efficiency factor $\kappa_{\mathrm{sw}}=1-\kappa$, we use the expression Lewicki:2022pdb ,
$\kappa=\mathcal{K}\dfrac{R_{\mathrm{eq}}}{R}\dfrac{\gamma}{\gamma_{\mathrm{eq}}},$ (29)
where $R$ is the radius of the bubble, $\gamma$ is the Lorentz factor of the bubble wall, $R_{\mathrm{eq}}$ is the radius of the bubble when it does not accelerate anymore, $\gamma_{\mathrm{eq}}$ is the Lorentz factor of the bubble wall when $R=R_{\mathrm{eq}}$ and finally $\mathcal{K}$ is a numerical prefactor which contains information of how much of the energy goes into the surrounding fluid and gets converted into thermal energy. For strongly supercooled FOPTs $\mathcal{K}\rightarrow 1$ and in general cases it can be expressed as,
$\mathcal{K}=\left(1-\dfrac{\alpha_{\infty}}{\alpha}\right)\left(1-\dfrac{1}{\gamma_{\mathrm{eq}}^{c}}\right),$ (30)
where $\alpha_{\infty}=\Delta P_{1\rightarrow 1}/\rho_{\mathrm{rad}}$ and $\Delta P_{1\rightarrow 1}$ denotes the pressure difference across the bubble wall due to one to one scattering events. Furthermore, in this study we consider $c=1$. For further insight on these quantities see Refs. Lewicki:2022pdb ; Ellis:2019oqb .
@@ 3.2 Creation of Primordial Black Holes
Now we discuss the creation of PBHs from a FOPT. The basic mechanism is that since the nucleation of true vacuum bubble is probabilistic process, there may exist some hubble region in which bubble nucleation has not taken place, i.e., that region is still in false vacuum and dominated by vacuum energy whereas its surrounding regions are in true vacuum and dominated by radiation. As a result, due to difference in the nucleation time, large inhomogeneities can generate which can in turn collapse to form PBH. In this regard, we follow the approach of Ref. Lewicki:2024ghw .
The horizon mass during the collapse of the overdense region can be expressed as,
$M_{H}=10^{32}\left(\dfrac{100}{g_{*}}\right)^{1/2}\left(\dfrac{T_{\mathrm{reh}}}{\mathrm{GeV}}\right)^{-2}~{}~{}\mathrm{g}.$ (31)
The abundance of the PBH created through this mechanism is given by Lewicki:2024ghw ,
$f_{\mathrm{PBH}}\approx 5.5\times 10^{6}\exp\left(-0.064e^{0.806\beta/H}\right)\left(\dfrac{g_{*}}{g_{*s}}\right)\left(\dfrac{T_{\mathrm{reh}}}{\mathrm{GeV}}\right).$ (32)
Furthermore, the mass function for these PBHs can be expressed as Lewicki:2024ghw ,
$\psi(M)\propto(M/M_{H})^{1+1/\gamma}\exp\left(-c_{1}(M/M_{H})^{c_{2}}\right),$ (33)
where $\gamma$ is the coefficient of the critical scaling law of the PBHs which takes the form $M=\kappa M_{k}(\delta-\delta_{c})^{\gamma}$ where $\kappa$ is the numerical prefactor, $M_{k}$ is the horizon mass when the perturbation enters the horizon, $\delta$ is the overdensity of the region and $\delta_{c}$ is the overdensity threshold beyond which a region collapses as the radiation pressure can not balance the overdensity. In this case, fixed values of $\gamma=0.36$, $\kappa=4$, and $\delta_{c}=0.5$ have been considered. It is also to be noted that the values $c_{1}$ and $c_{2}$ have weak dependence on $\beta/H$ and they take approximate values of $1.2$ and $2.7$ respectively for $\beta/H=8$ Lewicki:2024ghw . Therefore, it is trivial to see that in case of a FOPT of this type, the maximum PBHs of a certain population will have a certain mass $M_{\mathrm{PBH,~{}peak}}\approx 0.928M_{H}$, which we term as ‘peak mass’.
@@ 4.2 Parameter Scan
We can further divide the parameter scan effectively into two parts, namely, (i) $\chi$-related parameters, and (ii) $\Phi_{1,2}$-related parameters.
@@ 4.2.1 $\chi$-Related Parameters
In this part, we discuss the parameter ranges relevant for our work to understand the properties of the first FOPT, i.e. the one driven by the scalar field $\chi$. The parameters in question for this case are $g_{B-L}$, $y_{1,2,3}$, and $v_{\chi}$. As mentioned before, these parameters control the strength ($\alpha$), the inverse of the duration ($\beta/H$) and the relevant temperatures ($T_{p},~{}T_{n},~{}T_{c}$) of the FOPT, and those in turn control the PBH mass and abundance and GW spectrum. At first in Fig. 1 we show the dependence of $\alpha$ (left panel) and $\beta/H$ (right panel) on the couplings $g_{B-L}$ with varying Yukawa couplings while keeping the value of the VEV of the $\chi$ field fixed at $v_{\chi}=10^{8}\mathrm{~{}GeV}$.
@@
Figure 1: Dependence of (left) $\alpha$ and (right) $\beta/H$ values on $g_{B-L}$ for different values of $y_{1,2,3}$. In this case $v_{\chi}=10^{8}\mathrm{~{}GeV}$.
@@
It can be seen from Fig. 1 that for a fixed value of $y_{1,2,3}$, $\alpha$ ($\beta/H$) increases (decreases) significantly as $g_{B-L}$ decreases. On the other hand lower value of the Yukawa coupling leads to lower (higher) value of $\alpha$ ($\beta/H$). It is also to be noted that both $\alpha$ and $\beta/H$ depend very weakly on $y_{1,2,3}$ as compared to $g_{B-L}$. This behaviour is also true for the dependence $\alpha$ and $\beta/H$ on the $v_{\chi}$. We illustrate this further in Fig. 2 where we show the dependence of $\alpha$ (left panel) and $\beta/H$ (right panel) on $g_{B-L}$ for different values of VEV for constant Yukawa coupling $y_{1,2,3}=0.2$.
@@
Figure 2: Dependence of (left) $\alpha$ and (right) $\beta/H$ values on $g_{B-L}$ for different values of $v_{\chi}$. In this case $y_{1,2,3}=0.2$.
@@
It can be seen from the figure that for higher value of VEV, $\alpha$ is higher and $\beta/H$ is lower.
Figure 3: Dependence of $T_{n}$ on $g_{B-L}$ (left) and $y_{1,2,3}$ while keeping $v_{\chi}=10^{8}\mathrm{~{}GeV}$ (right) and $v_{\chi}$ keeping $y_{1,2,3}=0.2$.
@@
Apart from $\alpha$ and $\beta/H$, the temperatures related to the FOPT, i.e., $T_{p}$, $T_{n}$, and $T_{c}$ are also important since the peak of the SGWB and the mass of the PBH depends heavily on them. We found that for fixed VEV (Yukawa coupling), all three temperatures have similar behaviour with varying gauge and Yukawa coupling (VEV). Therefore in Fig. 3 we only show the $T_{n}$ dependence on $g_{B-L}$ and $y_{1,2,3}$ keeping $v_{\chi}=10^{8}\mathrm{~{}GeV}$ (left panel) and $T_{n}$ dependence on $g_{B-L}$ and $v_{\chi}$ keeping $y_{1,2,3}=0.2$ (right panel).
@@
From the figure, it can be seen that at fixed VEV and Yukawa coupling, $T_{n}$ increases with increasing $g_{B-L}$. In this case as well, $T_{n}$ weaker dependence on the Yukawa couplings as compared to $g_{B-L}$. However, for fixed gauge and Yukawa couplings, $T_{n}$ increases with VEV.